OpenStudy (anonymous):

Can someone check my answer? shayla is baking muffins which require 1/2 cup vegetable oil. she only has 1/3 cup vegetable oil left how much more does she need? my answer: 1/2-1/3 3/6-2/6=1/6 answer: 1/6

OpenStudy (anonymous):

@AskMeLotss sorry if i'm tagging you too much

OpenStudy (theyankee):

That's correct ^^

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

im here, sorry I took so long

OpenStudy (anonymous):

1/2 - 1/3 = 3/6 - 2/6 = 1/6 your answer is correct